# A Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor

### v2

@@ Abstract
The continuously parametrization in terms of $\ell$ can be exploited in different physical contexts. The systems we consider here could be realized in electronic systems by lithographic patterning of solid-state systems [140], depositing CO molecules on metals [141, 142], or by engineering 2D Shiba states with magnetic ad-atoms on superconductors [143, 144, 145]. Moiré heterostructures have also been used to realize unconventional tessellations [146]. Choosing different $\ell$ provides a new tuning knob to design topology and geometry in photonic metamaterials [147, 148, 6, 149, 150], polaritonic systems [151], electrical circuits [152], microwave networks [153], acoustic [154] and mechanical [155, 156, 157] metamaterials. More broadly, the tunability of this family of aperiodic tilings has already proven useful in enhancing earthquake detection [86] and could benefit other areas such the study of spin models [158], elasticity [159, 160], and fluids [161] on aperiodic tilings.
Acknowledgments — We thank E. Ben Achour, P. D’Ornellas, F. Flicker, L. Gómez Paz, Q. Marsal, S. Pachhal, M. Thumin, D. Varjas, and M. Yoshii for useful discussions and suggestions. We are grateful to P. D’Ornellas and S. Roche for their critical reading of the manuscript. J. S. is supported by the program QuanTEdu-France n° ANR-22-CMAS-0001 France 2030. A. G. G. is supported by the European Research Council (ERC) Consolidator grant under grant agreement No. 101042707 (TOPOMORPH).
Supplementary Material for “A Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor”
@@ Appendix A Generating aperiodic tilings from inflation rules
In this section, we describe the inflation procedure introduced in [66] and employed in [162] to generate the aperiodic tilings we use in the main text. As illustrated in Fig. S1, this procedure involves the successive replacement of one set of tiles with another made of a larger amount of tiles. In this work, we adapt the code used in Ref. [86] to generate the tilings presented in Figs. 1 and 2. We begin by examining a single tile, which is a 14-sided polygon whose edges can be grouped into pairs of equal-length, parallel segments, as shown in Fig. S1a. The first step is to combine one tile with its mirror image, forming a two-tile compound. Following [66], this two-tile compound can then be combined with six additional tiles to create a cluster, referred to as H8, or with five other tiles to form the H7 cluster. The inflation process proceeds as follows: each individual tile is replaced by an H8 cluster, while each two-tile compound is replaced by an H7 cluster. This process can be iterated indefinitely to generate arbitrarily large lattices. The number of tiles after the $n$-th iteration is given by the following recursion relation
@@
where $\mathcal{H}$ is the real-space Hamiltonian and $E$ is the energy at which the Green’s function is computed. In practice, $\eta$ is taken to be finite. The value of $\eta$ is related to the number of moments in the expansion in terms of Tchebychev’s polynomials used in the kernel polynomial method [83]. We now introduce the tight-binding basis $\lvert{\bf{r}_{\alpha}},l\rangle$ that span the Hilbert-space. Each $\lvert{\bf{r}_{\alpha}},l\rangle$ is localized on a site $\alpha$ and carries an internal orbital degree of freedom labeled $l$. These states are represented as $N_{\text{Sites}}\times N_{\text{Orbs}}$ vectors and satisfy the following relation
@@ Appendix D Spectral localizer gap and choice of $\kappa$
Figure S3: Localizer gap of the spectral localizer Eq. (19) as a function of the reference energy $E_{0}/t$ and the parameter $\kappa$, measured in units of $t/\ell_{c}$. Here $\ell_{c}=1$ is the side of the Comet tile, see Fig. 1a, and sets our unit length. We compute in the bulk of the system, at a position $x_{0},y_{0}$ chosen by computing the geometric mean of the site’s coordinates, for $\ell=0.5$ and $M/t=0.5$. The vertical dashed black line corresponds to the value of $\kappa$ used in the main text. For $E_{0}/t=2.65$, the energy set by $E_{F}$ at this $\ell$ in Fig. 3, we are within a bubble with a large localizer gap that is well defined for several orders of magnitude in $\kappa$.
@@
In this section, we justify the choice of the parameter $\kappa$ used to generate Fig. 3 via the spectral localizer [103, 104, 105, 106], whose definition we repeat here for convenience
=(^$\mathcal{L}=(\hat{\mathcal{H}}-E_{0}\mathbb{1})\sigma_{z}+\kappa\left[(\hat{X}-x_{0}\mathbb{1})\sigma_{x}+(\hat{Y}-y_{0}\mathbb{1})\sigma_{y}\right].$ (19)
The parameter $\kappa$ balances the typical energy scales of the Hamiltonian $\mathcal{H}$ compared to those of the the position operators $\hat{X}$ and $\hat{Y}$. If $\kappa$ is chosen too small, the localizer $\mathcal{L}$ becomes approximately (^$\left(\hat{\mathcal{H}}-E_{0}\mathbb{1}\right)\sigma_{z}$, whose spectrum is symmetric around zero. Because of this symmetry the signature vanishes, yielding a trivial topological invariant. Conversely, if $\kappa$ is chosen too large, the localizer is dominated by the position operators and effectively decouples from the Hamiltonian, again resulting in a trivial invariant. Therefore, $\kappa$ must be chosen so that the Hamiltonian and the position operators contribute comparably to the localizer, ensuring the computation captures the correct topological features.
To find a suitable value of $\kappa$, it is convenient to compute the localizer gap at different energies for different values of $\kappa$ [106]. The localizer gap is defined as the magnitude of the smallest eigenvalue in absolute value, $\min(\lvert\lambda\rvert)$ where $\lambda\in\mathrm{Sp}(\mathcal{L})$. For a given matrix $A$, $\mathrm{Sp}(A)$ denotes the set of all the eigenvalues of $A$ or spectrum of $A$. Because the invariant can only change when the localizer gap closes, the magnitude of the localizer gap attests to the robustness of a given phase computed for a given choice of parameters $(\kappa,x_{0},y_{0},E_{0})$. A large gap, extended in parameter space, signifies a stable phase [105, 167, 168, 106]. In this sense the localizer gap acts analogously to the band-gap of crystalline systems. In practice, for a given system size there is not a single value of $\kappa$ but rather a window of values of $\kappa$ that capture well the topologically robust properties of the system [106].
@@
In Fig. S3 we plot the localizer gap as a function of both $\kappa$ and $E_{0}/t$ for $x_{0}$ and $y_{0}$ chosen in the bulk of the system, by computing the geometric mean of the site’s coordinates. We observe the emergence of distinct bubbles where the localizer gap is large, similar to other localizer studies [105, 167, 168, 106, 169]. These bubbles correspond to the regions in parameter space where both the Hamiltonian and the position operators enter meaningfully the calculation of the spectral localizer. We have checked that, by selecting a value of $\kappa$ such that the energy of interest lies within one of these bubbles, we ensure that the invariant computed via the spectral localiser at a given energy approximately coincides with the bulk average of the local Chern marker at the same energy. Furthermore, we observe that the bubble at $E_{0}/t=2.65$, the energy set by $E_{F}$ at this $\ell$ in Fig. 3, is particularly extended, which ensures the robustness of the calculation with respect to different choices of the parameter $\kappa$.

### v4

@@ Abstract
The continuous parametrization in terms of $\ell$ can be exploited in different physical contexts. The systems we consider here could be realized in electronic systems by lithographic patterning of solid-state systems [146], depositing CO molecules on metals [147, 148], or by engineering 2D Shiba states with magnetic ad-atoms on superconductors [149, 150, 151]. Moiré heterostructures have also been used to realize unconventional tessellations [152]. Choosing different $\ell$ provides a new tuning knob to design topology and geometry in photonic metamaterials [153, 154, 10, 155, 156, 157], polaritonic systems [158], electrical circuits [159], microwave networks [160], acoustic [161] and mechanical [162, 163, 164] metamaterials. More broadly, the tunability of this family of aperiodic tilings has already proven useful in enhancing earthquake detection [89] and could benefit other areas such as the study of spin models [165], elasticity [166, 167], and fluids [168] on aperiodic tilings.
Acknowledgments — We thank E. Ben Achour, P. D’Ornellas, F. Flicker, L. Gómez Paz, Q. Marsal, S. Pachhal, M. Thumin, D. Varjas, and M. Yoshii for useful discussions and suggestions. We are grateful to P. D’Ornellas and S. Roche for their critical reading of the manuscript. Technical and human support provided by the DIPC Supercomputing Center is gratefully acknowledged. J. S. is supported by the program QuanTEdu-France n° ANR-22-CMAS-0001 France 2030. A. G. G. is supported by the European Research Council (ERC) Consolidator grant under Grant Agreement No. 101042707 (TOPOMORPH).
The data and codes that support the findings of this article are openly available at [169].
Supplementary Material for “Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor”
@@ Appendix A Generating aperiodic tilings from inflation rules
In this Appendix, we describe the inflation procedure introduced in [70] and employed in [170] to generate the aperiodic tilings we use in the main text. As illustrated in Fig. S1, this procedure involves the successive replacement of one set of tiles with another made of a larger amount of tiles. In this work, we adapt the code used in Ref. [89] to generate the tilings presented in Figs. 1 and 2. We begin by examining a single tile, which is a 14-sided polygon whose edges can be grouped into pairs of equal-length, parallel segments, as shown in Fig. S1a. The first step is to combine one tile with its mirror image, forming a two-tile compound. Following [70], this two-tile compound can then be combined with six additional tiles to create a cluster, referred to as H8, or with five other tiles to form the H7 cluster. The inflation process proceeds as follows: each individual tile is replaced by an H8 cluster, while each two-tile compound is replaced by an H7 cluster. This process can be iterated indefinitely to generate arbitrarily large lattices. The number of tiles after the $n$-th iteration is given by the following recursion relation
@@
where $\mathcal{H}$ is the real-space Hamiltonian and $E$ is the energy at which the Green’s function is computed. In practice, $\eta$ is taken to be finite. The value of $\eta$ is related to the number of moments in the expansion in terms of Chebychev polynomials used in the kernel polynomial method [86]. We now introduce the tight-binding basis $\lvert{\bf{r}_{\alpha}},l\rangle$ that span the Hilbert-space. Each $\lvert{\bf{r}_{\alpha}},l\rangle$ is localized on a site $\alpha$ and carries an internal orbital degree of freedom labeled $l$. These states are represented as $N_{\text{Sites}}\times N_{\text{Orbs}}$ vectors and satisfy the following relation
@@ Appendix D Spectral localizer gap and choice of $\kappa$
Figure S3: Localizer gap of the spectral localizer Eq. (19) as a function of the reference energy $E_{0}/t$ and the parameter $\kappa$, measured in units of $t/\ell_{c}$, using a logarithmic colormap. Here $\ell_{c}=1$ is the side of the Comet tile, see Fig. 1a, and sets our unit length. We compute in the bulk of the system, at a position $x_{0},y_{0}$ chosen by computing the geometric mean of the site’s coordinates, for $\ell=0.5$ and $M/t=0.5$. The vertical dashed black line corresponds to the value of $\kappa$ used in the main text. For $E_{0}/t=2.65$, the energy set by $E_{F}$ at this $\ell$ in Fig. 3, we are within a bubble with a large localizer gap that is well defined for several orders of magnitude in $\kappa$.
@@
In this Appendix, we justify the choice of the parameter $\kappa$ used to generate Fig. 3 via the spectral localizer [106, 107, 108, 109], whose definition we repeat here for convenience
$\mathcal{L}=(\hat{\mathcal{H}}-E_{0}\mathbb{1})\sigma_{z}+\kappa\left[(\hat{X}-x_{0}\mathbb{1})\sigma_{x}+(\hat{Y}-y_{0}\mathbb{1})\sigma_{y}\right].$ (19)
[…]
@@
In Fig. S3 we plot the localizer gap as a function of both $\kappa$ and $E_{0}/t$ for $x_{0}$ and $y_{0}$ chosen in the bulk of the system, by computing the geometric mean of the site’s coordinates. We observe the emergence of distinct bubbles where the localizer gap is large, similar to other localizer studies [108, 175, 176, 109, 177]. These bubbles correspond to the regions in parameter space where both the Hamiltonian and the position operators meaningfully enter the calculation of the spectral localizer. We checked that, by choosing a value of $\kappa$ such that the energy of interest falls within one of these bubbles, the invariant obtained from the spectral localizer at that energy closely matches the bulk average of the local Chern marker at the same energy, see next section for a more detailed discussion.
Furthermore, we observe that the bubble at $E_{0}/t=2.65$, the energy set by $E_{F}$ at this $\ell$ in Fig. 3, is particularly extended, which ensures the robustness of the calculation with respect to different choices of the parameter $\kappa$.
Appendix E Comparison between the local Chern marker and the spectral localizer index
Formally, the spectral localizer index and the local Chern marker coincide in the small $\kappa$ limit [111]. Small $\kappa$ is defined with respect to the ratio between the typical energy scales of the Hamiltonian, e.g. the band width, and the linear system size $L$. Hence, for a given $L$ and Hamiltonian parameters, we can find a small $\kappa$ that makes the localizer index and average local Chern marker in the bulk approximately equal. In practice, we follow the convention in the literature to fix a single $\kappa$ to calculate our phase diagrams [109]. This is similar to the usual convention of fixing the region of integration of the local Chern marker, despite the fact that its bulk average changes depending on the spatial extent of the edge states, governed by the Hamiltonian parameters.
With these caveats in mind we now show that, in practice, fixing $\kappa$ for the full phase diagram recovers compatible numerical results between both methods in a large phase space, and that the agreement can be made increasingly better by tuning $\kappa$ [111]. In Fig. S4a-b we compare the phase diagrams of the bulk average of the local Chern marker, $C$, and the local quantum metric, $G$, and the localizer index for two different values of $\kappa$ ($\kappa=0.01,0.001$). Fig. S4 highlights two main features. First, the phase diagrams of the localizer index and local Chern marker become increasingly similar as $\kappa$ is reduced, see Fig. S4a. Second, the deviations occur at regions where the bulk average of the quantum metric is large, see Fig. S4b. This indicates that the discrepant regions have more extended wavefunctions due to smaller spectral gaps. These properties are also reflected in the small size of the localizer gap, as seen in Fig. 3b. By reducing $\kappa$ the spectral localizer gains energy resolution, per Eq. (19), and is thus able to resolve topological properties of finer energy gaps, and the two phase diagrams become increasingly similar, as predicted by Ref. [111]. Hence, we conclude that for a given system size, smaller $\kappa$ enlarges the nontrivial regions identified by the localizer and brings the two methods into closer correspondence.
We stress that, independent of the choices above, the phase diagram shows topological regions that are larger than for the square-lattice crystalline model, as described in the main text. Throughout the main text, we have fixed $\kappa=0.01$ for the system sizes considered. As can be seen in Fig. S4a this choice is conservative.
Figure S4: Topological phase diagrams and quantum metric. (a) Phase diagram obtained from the bulk average of the local Chern marker, $C=(1/\mathcal{A}_{b})\sum_{\alpha\in\text{Bulk}}C_{\textbf{r}_{\alpha}}$, see Eq. (2), for a system of 2530 sites. The normalization area $\mathcal{A}_{b}$ is defined as in Fig. S2. The phase boundaries of the spectral localizer index are shown in black for $\kappa=0.01$ (dashed) and $\kappa=0.001$ (solid). (b) Phase diagram obtained from the bulk trace of the local quantum metric $G$ for the same system, see Eq. (18). Smaller values of $\kappa$ bring the phase diagrams computed via the spectral localizer and local Chern marker closer to agreement by capturing the topology of regions with larger quantum metric and smaller bulk gaps.
Appendix F Finite-size and disorder-averaging effects
Figure S5: Convergence of the phase diagram in the presence of Anderson disorder at 3/4 filling. (a) Half-signature of the spectral localizer as a function of disorder strength $W/t$ averaged over 100 disorder realizations for three different system sizes with 404, 2530 and 16496 sites, generated by two, three and four iterations of the meta-tile construction, respectively. Shaded regions indicating the standard error. (b) Disorder-averaged half-signature at $\ell=0.5$ for different number of disorder realizations, with standard error shown as shaded regions. (c) and (d) show the phase diagram obtained from 50 and 100 disorder realizations, respectively.
In this Appendix, we examine the effect of increasing the number of disorder realizations averaged on the results presented in the main text and demonstrate that our results are converged with respect to system size and disorder realizations. Fig. S5a shows the half-signature at fixed parameters $\ell=0.6$ and M/t=2.7M/t=-2.7, where we vary the disorder strength $W/t$ for three system sizes: a small, medium and large system with 404, 2530 and 16496 sites generated by two, three and four iterations of the meta-tile construction, respectively. All sizes lead to a similar profile of the topological transition as a function of disorder strength. We note that the width of the transition region, where the marker deviates from integer values, is relatively broad but comparable to those reported for a quantum spin-Hall Hamiltonian in the Ammann-Beenker tiling [178]. The model used in Ref. [178] is close to a doubled version of our Hamiltonian, with a copy per spin. Fully characterizing this transition is an interesting future research direction.
In Fig. S5b, we present a cut at fixed $\ell=0.5$, averaged over different numbers of disorder realizations, ranging from 50 to 500, for a system with 2530 sites. The results converge already around 100 realizations, with only negligible differences between 100 and 500. To further confirm this convergence, Figs. S5c and d display the $(W/t,\ell)$ phase diagram for a system with 2530 sites, averaged over 50 and 100 disorder realizations, respectively. The two phase diagrams exhibit nearly identical phase boundaries and regions, demonstrating convergence.
Appendix G Example of the local quantum metric in a topological and trivial phases
In this Appendix, we illustrate how the local quantum metric is spatially distributed in both trivial and topological phases. Figure S6 shows the quantum metric computed at M/t=2.7M/t=-2.7 and 3/4 filling for a system with 2530 sites. With $\ell=0.1$ and $\ell=0.5$ the system lies in the trivial and topological phases, respectively. In the trivial phase, the quantum metric is localized within small clusters of sites, see Fig. S6a. This is strongly indicative of localized wavefunctions, consistent with the state being in a trivial phase. By contrast, in the topological phase, the local quantum metric is nearly uniform in the bulk and reaches its maximum along the system’s edge, see Fig. S6b. This is consistent with presence of a topological bulk with delocalized topological edge states.
Figure S6: Real-space distribution of the local quantum metric, $g_{\mathbf{r}}$. (a) Local quantum metric computed using Eq. (17) in the trivial phase at 3/4 filling with M/t=2.7M/t=-2.7 and $\ell=0.1$. (b) Same calculation for $\ell=0.5$ in the topological phase. In the trivial case (a), the metric is concentrated in localized clusters, whereas in the topological case (b), it is nearly uniform in the bulk and enhanced at the edges, reflecting the presence of a delocalized topological edge state.